given a model with the following structural quations
y = c + I + G + (X - M) where
C = 30 + 0.75Y
I = N80 million
G = 120 million
X = 20 million
M = 5 + 0.2 y
determine
i) equilibrium national income
ii) income multiplier
b analyze the import demand equation
use a numerical example to define and explain the concept of multipier
1
Expert's answer
2020-12-28T08:27:06-0500

"Y=30+0.75Y+80+120+(20-5+0.2Y)"

"Y=30+80+120+20-5+0.75Y+0.2Y"

"Y=110+120+15+0.95Y"

"Y=245+0.95Y"

"Y-0.95Y=245"

"0.05Y=245"

"Y= \frac{245}{0.05}"

"Y=4900"

Income Multiplier="\frac{Y}{G}"

"= \frac{4900}{120}"

"=40.8"

"=41"

Import Demand Equation

"M=5+0.2Y"

"=5+0.2(4900)"

"=5+980"

"=985"

Multiplier is the link amid the government expenditure and total national income. For instance, 2x would double the base figure while 0.5x would reduce the base figure by half.